OpenStudy (anonymous):

"1/R effective = 1/12ohms + 1/15ohms" And just so you know, it's like the ohms symbol is in the denominator with the numbers, not to the side of it. Could you please solve and explain this problem?

OpenStudy (anonymous):

This equation is used when resistors are in parallel. if R1 and R2 are two resistors in parallel and the total resistance Rt of the combination, then 1/Rt = 1/R1 + 1/R2. As per your equation you a 12ohm resistor and a 15ohm resistor in parallel. 1/Reffective = 27/12*15 = 3/20 therefore Reffective = 20/3

OpenStudy (anonymous):

Take the reciprocal of both sides ad simplify by hand or evaluate on a calculator. Resultant units will be in ohms as well.\[R_{effective}=\frac{1}{\dfrac{1}{15\ \Omega}+\dfrac{1}{12\ \Omega}}\]

OpenStudy (anonymous):

take the reciprocal after simplifying.